A 4 kg stone is whirled at a constant speed in a vertical circle of radius 1 m. At the bottom of the circle the tension is 2mg. The speed of the stone is _____ m/s?
Answer #1

at bottom

T = mv^2/r + mg

2mg = mv^2/r + mg

2g = v^2/r + g

v^2/r = g

v = sqrt(rg)

v = 3.13 m/s